OpenStudy (anonymous):

Write a rule for a quadratic function that has x-intercepts of -3 and 4 with a y-intercept if 10. Show all work.

OpenStudy (anonymous):

you know it looks like \[f(x)=a(x+3)(x-4)\] and you know \[f(0)=10\] so you can solve for "a" via \[a\times 3\times -4=10\]

OpenStudy (anonymous):

x-intercept => y = 0, x-intercept1 = -3 and x- intercept2 =4, means then that when y = 0, x = -3, we have (-3,0) or when y= 0, x = 4, (4,0) y-intercept => x =0 (0,10) y= ax^2 + bx + c =0, x= 4, x = -3 are roots, or solutions. When x = 0, y= a * 0^2 + b * 0 +c, y = c. When x is 0, y is 10, then c = 10. a(-3)^2 -3b + 10 = 0 9a -3 b = - 10 a (4)^2 + 4b + 10 = 0 16 a + 4 b = - 10

OpenStudy (anonymous):

9a - 3b = - 10 16 a + 4b = - 10 9a - 3 b =16 a + 4b - 7 a = 7 b - a = b

OpenStudy (anonymous):

ax^2 -ax + 10 = 0 â (x^2 - x) = - 10 a(16 - 4) = - 10 a * (12) = - 10 a = -5/6 b = 5/6 -5/6 x^2 + 5/6 x + 10 = 0

OpenStudy (anonymous):

wow i think that is right but it is a lot of work. you can write \[f(x)=a(x+3)(x-4)\] and \[10=f(0)=a\times 3\times -4\] \[-12a=10\] \[a=-\frac{5}{6}\] so your answer is \[f(x)=-\frac{5}{6}(x+3)(x-4)\] multiply out if you like